OpenStudy (anonymous):

Which is the same as 4 m? 0.04cm 400cm 4000km 400km

OpenStudy (anonymous):

400 cm

OpenStudy (anonymous):

1m = 100cm 4m = 400cm

OpenStudy (angela210793):

1m=100cm 4m=xcm x=4*100=400cm

OpenStudy (anonymous):

\[4 meters \frac{100 centimeters}{1 meters} = 400 centimeters\]

OpenStudy (anonymous):

1m=100cm 4m=xcm x=4*100=400cm